### v2

# Behavior-Inspired Neural Networks for Relational Inference

Abstract
From pedestrians to Kuramoto oscillators, interactions between agents govern how a multitude of dynamical systems evolve in space and time. Discovering how these agents relate to each other can improve our understanding of the often complex dynamics that underlie these systems. Recent works learn to categorize relationships between agents based on observations of their physical behavior. These approaches are limited in that the relationship categories are modelled as outcomes of categorical distribution, when in real world systems categories often intermingle and interact. In this work, we introduce a level of abstraction between the observable behavior of agents and the latent categories that determine their behavior. To do this, we learn a mapping from agent behavior to agent preferences for each latent category in a graph neural network. We integrate the physical proximity of agents and their preferences in a nonlinear opinion dynamics model which provides a mechanism to identify mutually exclusive latent categories, predict an agent’s evolution in time, and control an agent’s physical behavior. We demonstrate the utility of our model for learning interpretable categories, and its efficacy on long-horizon prediction across several benchmarks where we outperform existing methods.
1 Introduction
Figure 1: Behavior-inspired neural network for relational inference. BINNs allow for relational inference using inter-agent distance and a learned representation of inter-category interactions. The physical states of agents are encoded to preferences for a set of latent categories. Preferences are propagated forward in time using a nonlinear opinion dynamics model with learned parameters. Predicted preferences are then decoded to physical states as predicted future physical states.
Multi-agent systems are found in domains as diverse as astronomy [1, 2], biology [3, 4, 5], physics [6, 7], and sports [8, 9]. Understanding how these systems evolve in time has the potential to provide insights useful for the discovery of unknown physics, the rules governing collective behavior, and engineering design. Predicting the evolution of complex systems is a fundamental challenge in the learning literature. Early black-box models determine future states from past states without regard for contextual information [10, 11]. More recent approaches improve upon these models by incorporating information about environmental conditions and the behavior of other agents [12, 13, 14, 15, 16]. While incorporating contextual information has led to more powerful trajectory prediction models, the opacity of these models limits our ability to leverage them to better understand the role of inter-agent relationships.
Recent work in relational inference attempts to address this limitation by explicitly modelling inter-agent relationships [17, 18, 19, 20, 21, 22]. Discovering inter-agent relationships is challenging, however, since ground truth labels are typically unavailable, and the relevant relationships may be unknown at design time. Graph neural network (GNN) based approaches such as  [19, 20, 21, 22] model inter-agent relationship categories with a categorical variable. In their seminal work, the authors of [19] learn a latent representation of inter-agent relationships in a trajectory prediction pipeline. The authors of [20] improve on this approach by modeling inter-agent relationships as time varying and the authors of [21] improve model expressivity by representing node features as a hypergraph. While these methods are able to predict future observations of a number of systems, the underlying assumption that inter-agent relationships are determined by a single relationship category diverges from what we observe in the real world.
In contrast to these models, opinion dynamics models assume agents have preferences for multiple categories, and that these preferences evolve in time. The nonlinear opinion dynamics (NOD) model introduced in [23], has been used to model a variety of societal systems [24, 25, 26, 27, 28, 29, 30, 31]. The nonlinear nature of this model introduces a bifurcation which allows preferences to quickly and flexibly respond to changes in environmental inputs. A limitation of this method, however, is that the relationship between agent preferences for a set of categories and their observable behavior must be known a priori.
In our model, Behavior-inspired Neural Network (BINN), we combine the flexibility of GNN models with the interpretability of nonlinear opinion dynamics for a new approach to relational inference. Concretely, our contributions are the following:
  • In contrast to existing opinion dynamics models which require observability of agent preferences, we learn a representation of agent preferences from physical observations.
  • In contrast to existing relational inference approaches which model relationship categories as a categorical variable, we use the NOD inductive bias to model the evolution of real-valued agent preferences on a set of latent categories.
  • By incorporating the NOD inductive bias, we can identify mutually exclusive categories, predict an agent’s evolution in time, and control an agent’s behavior.
We demonstrate the utility of our approach for identifying mutually exclusive categories on multiple illustrative examples and demonstrate the efficacy of our approach for trajectory prediction tasks.
2 Related Work
Trajectory prediction.
In contrast to traditional control approaches which tune a set of model parameters from all available trajectories (i.e., sequences of physical positions) [32, 33, 34, 35], early deep learning approaches learn to map a sequence of input states to future states directly [11]. Later efforts incorporated external influences (e.g., behavior of other agents, environmental conditions) [12, 13, 14, 15, 16] and temporal dependencies [16, 36, 37, 38, 39], and recent work has used graph based methods to model multi-agent dynamics [40, 41]. While these methods are able to predict future system states , they do not predict which agents interact or how they interact, a limitation which motivates relational inference.
Relational inference.
The goal of relational inference is to infer inter-agent relationships in a multi-agent system [17, 19, 20, 21, 22]. This task is challenging since, in general, the relationships between agents is unknown. Approaches to this challenge include: CommNet [17] which learns a continuous communication protocol for fully cooperative multi-agent tasks; Neural Relational Inference [19] which learns time invariant inter-agent relationships; Dynamic Neural Relational Inference [20] which learns time varying inter-agent relationships; GroupNet [21] which uses a multiscale hypergraph representation of the input for improved expressivity; and EquiMotion [22] which incorporates an equivariance constraint to ensure the dynamics transform predictably to Euclidean shifts of the input data. Our work differs from these approaches in our representation of relationship categories. Our categories exist in the space of agent preferences rather than the physical space, and our categories are flexible and interacting, rather than mutually exclusive and independent.
Consensus dynamics.
In control and robotics, consensus dynamics [42] have been used in a myriad of settings to model the dynamics and control the behavior of multi-agent systems. In [43], the authors propose a linear model for prediction and control of multi-agent systems. In [44], the authors use a linear consensus dynamics model for coordinated surveying with underwater gliders. In [45], the authors use the consensus dynamics framework to develop rectilinear and circular consensus control laws for multi-vehicles systems. In [46], the authors use consensus dynamics to improve data collection in a mobile sensor network. In [47], the authors use consensus dynamics to understand the robustness of starling flocking behavior. Even with this breadth of application, there are several draw backs to a linear model of opinion formation; specifically, the naive implementation results in dynamics that only yields consensus [48, 49] and the formation of opinions in response to inputs is slow.
Nonlinear opinion dynamics.
The noted short comings of linear consensus dynamics models are resolved in the nonlinear opinion dynamics model proposed in [23]. As the model is nonlinear, the dynamics result in a bifurcation and opinions can evolve to dissensus quickly. Nonlinear opinion dynamics have been used to model a variety of systems. In [24, 25, 26, 27] nonlinear opinion dynamics model information spread in settings such as political polarisation. In [28], nonlinear opinion dynamics is used to resolve deadlock, and in [29], it is used for collision avoidance. In [30, 31], nonlinear opinion dynamics is used for task switching in robotic swarms in trash collecting robot teams. In contrast to these works, where the model has direct access agent opinions (i.e., preferences), we learn a mapping between physical states and agent preferences.
3 Background
Nonlinear dynamics differ from linear dynamics in that they are able to exhibit bifurcations [50, 51, 52]. Nonlinear opinion dynamics [53] leverage bifurcations for fast and flexible decision making even with weak input signals [23]. Our BINN model integrates a nonlinear opinion dynamics inductive bias to interpretably and flexibly model agent behavior.
3.1 Nonlinear opinion dynamics
Consider a multi-agent system of $\mathcal{N}_{\mathrm{a}}\in\mathbb{N}$ agents, each possessing real-valued preferences about $\mathcal{N}_{\mathrm{o}}\in\mathbb{N}$ categories. Each category represents a belief or desire (e.g., take a physical action, undertake a task), and an agent’s preference for a category can be positive, neutral, or negative (the preference magnitude corresponds to preference strength). Changes to an agent’s preference for a category depend on extrinsic and intrinsic parameters. Concretely, the changes to agent $i$’s preference for category $j$ can be determined by the following nonlinear differential equation proposed in [23]
{dmath}˙z
_ij = -d_ijz_ij + S(u_i(α_ijz_ij + ∑_k=1k≠i^N_aa^a_ikz_kj + ∑_l=1l≠j^N_oa^o_jlz_il + ∑_k=1k≠i^N_a∑_l=1l≠j^N_o a^a_ika^o_jlz_kl)) + b_ij.
The parameters $d_{ij}\geq 0$, $u_{i}\geq 0$, and $\alpha_{ij}\geq 0$ are intrinsic to the agent. The damping $d_{ij}$ represents how resistant agent $i$ is to forming a preference for category $j$, the attention $u_{i}$ represents how attentive agent $i$ is to the preferences of other agents, and the self-reinforcement $\alpha_{ij}$ represents how resistant agent $i$ is to changing its preference about category $j$.
The parameters $a^{a}_{jk}$, $a^{o}_{jl}$, and $b_{ij}$ are extrinsic to the agent. The communication matrix $[a^{a}_{ik}]$ describes the communication strength between agent $i$ and agent $k$, the belief matrix $[a^{o}_{jl}]$ describes the correlation of preferences for category $j$ and category $l$, and the environmental input $b_{ij}$ describes the impact of the environment on an agent’s the preference for category $j$. The saturating function $\mathcal{S}$ is selected so that $S\left(0\right)=0$, $S^{\prime}\left(0\right)=1$, and $S^{\prime\prime\prime}\left(0\right)\neq 0$ [23]. We use a hyperbolic tangent for the saturation function in our experiments.
For a given communication matrix $[a^{o}_{jl}]$ and belief matrix $[a^{o}_{jl}]$, the sensitivity of preference formation is determined by an agent’s intrinsic parameters $d_{ij}$, $u_{i}$, and $\alpha_{ij}$; and the equilibrium value of each preference is a function of the environmental input $b_{ij}$ (see Figure 2).
[…]
Figure 2: Sensitivity to environmental inputs $b_{ij}$. Solid lines represent stable equilibria and dotted line represents unstable equilibria. (a) We show the pitchfork bifurcation characteristic to equation 3.1. The number, location, and stability of equilibria changes with the attention parameter $u_{i}$. (b) For attention $u_{i}<u^{*}$ preferences change linearly with environmental inputs. (c) For attention $u_{i}>u^{*}$ preferences change rapidly in response to environmental inputs with hysteresis encoding memory of previous states.
[…]
Figure 3: Behavior-inspired neural network architecture overview. Our network takes agent states as inputs and outputs predicted next states, while maintaining a representation of agent preferences for a set of latent categories. The network $E_{z}$ encodes physical states to preferences for a set of latent categories and $E_{b}$ encodes physical states into latent environmental inputs. In the latent space, we compute future preferences using the nonlinear opinion dynamics block $f_{\mathrm{NOD}}$, and use the decoder $D_{x}$ to map predicted preferences to predicted physical states. The latent dynamics are unrolled for multi-step trajectory prediction.
Mutually exclusive categories.
In the nonlinear opinion dynamics framework, mutually exclusive categories are negatively correlated. Consider, for example, a system with two categories. The categories are mutually exclusive if $a^{o}_{12},a^{o}_{21}\leq 0$. Since $z_{i1}=-cz_{i2}$, we can decouple the dynamics for $z_{i1}$ and $z_{i2}$ and only model the preference for one category. Equation (3.1) therefore reduces to
$\dot{z}_{i}=-d_{i}z_{i}+\mathcal{S}\left(u_{i}\left(\tilde{\alpha}_{i}z_{i}+\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}\tilde{a}_{ik}z_{kj}\right)\right)+b_{i}.$
Additional details are provided in Appendix C.1. Mutually exclusive categories provides a systematic way to identify opportunities for dimensionality reduction.
4 Method
In this section we present our BINN model for relational inference (see Figure 3). Given trajectories of a multi-agent system, our goal is to predict agent behavior in an interpretable way by inferring the intrinsic and extrinsic characteristics of agents as described by the nonlinear opinion dynamics model (see Section 3).
We define a multi-agent system as a system of $\mathcal{N}_{\mathrm{a}}\in\mathbb{N}$ agents, each with real-valued preference for a set of $\mathcal{N}_{\mathrm{o}}\in\mathbb{N}$ categories. We learn a mapping between the observed behavior of agents and their preferences for a set of latent categories using a graph neural network; and the parameters of a nonlinear opinion dynamics model which determine the evolution of agent preferences.
We model the evolution of agent preference by the dynamical equation presented in Equation 3.1. With this formulation, we can control agent preferences by varying the environmental input as shown in Figure 5.
We train our model using $N$ trajectories of $T$ observations, where each observation has dimension $d$. For a given trajectory, we denote the observed state of agent $i$ at time $t$ by $\mathbf{x}_{i,t}\in\mathbb{R}^{d}$, and the preferences of agent $i$ at time $t$ by $\mathbf{z}_{i,t}\in\mathbb{R}^{\mathcal{N}_{o}}$.
4.1 Encoder
We use separate encoding networks to learn the mappings from agent states to agent preferences and from agent states to environmental inputs. We refer to the first of these encoders as the preference encoder $E_{z}$, and the latter as the environmental input encoder $E_{b}$. Each encoder takes state observations of the multi-agent system at time $t$ as input and processes them in an MPNN [54]. The multi-agent system is represented as a fully-connected graph with node values determined by the physical state of each agent.
Formally, the preference encoder $E_{z}$ performs the following message passing functions for agent $i$ at timestep $t$:
$\quad\quad\mathbf{z}^{\prime}_{i,t}$ $=f_{\mathrm{emb}}^{{z}}\left(\mathbf{x}_{i,t}\right),$ (2)
$v\rightarrow e:$ $\quad\quad\mathbf{m}^{z}_{\left(i,k\right),t}$ $=f_{v\rightarrow e}^{{z}}\left(\mathbf{z}^{\prime}_{i,t},\mathbf{z}^{\prime}_{k,t}\right),$ (3)
$e\rightarrow v:$ $\quad\quad\mathbf{z}_{i,t}$ $=f_{e\rightarrow v}^{{z}}\left(\textstyle\sum_{k\neq i}\mathbf{m}^{z}_{\left(i,k\right),t},\mathbf{z}^{\prime}_{i,t}\right),$ (4)
where $f_{\mathrm{emb}}^{z}$, $f_{v\rightarrow e}^{z}$, and $f_{e\rightarrow v}^{z}$ are 3-layer MLPs. The environmental input encoder $E_{b}$ is designed similarly, and performs the following message passing functions for agent $i$ at timestep $t$:
$\quad\quad\mathbf{b}^{\prime}_{i,t}$ $=f_{\mathrm{emb}}^{{b}}\left(\mathbf{x}_{i,t}\right),$ (5)
$v\rightarrow e:$ $\quad\quad\mathbf{m}^{b}_{\left(i,k\right),t}$ $=f_{v\rightarrow e}^{{b}}\left(\mathbf{b}^{\prime}_{i,t},\mathbf{b}^{\prime}_{k,t}\right),$ (6)
$e\rightarrow v:$ $\quad\quad\mathbf{b}_{i,t}$ $=f_{e\rightarrow v}^{{b}}\left(\textstyle\sum_{k\neq i}\mathbf{m}^{b}_{\left(i,k\right),t},\mathbf{b}^{\prime}_{i,t}\right),$ (7)
where $f_{\mathrm{emb}}^{{b}}$, $f_{v\rightarrow e}^{b}$, and $f_{e\rightarrow v}^{b}$ are 3-layer MLPs.
4.2 Latent nonlinear opinion dynamics
We use the nonlinear opinion dynamics formulation in Equation (3.1) to model the evolution of agent preferences on a set of latent categories,
$\dot{\mathbf{z}}_{i,t+1}=f_{\mathrm{NOD}}\left(\mathbf{z}_{i,t},\mathbf{b}_{t},\mathbf{A}^{\mathrm{a}}_{t}\right).$ (8)
We learn the intrinsic agent parameters, $\mathbf{d}$, $\mathbf{u}$, and $\boldsymbol{\alpha}$, the extrinsic parameter $\mathbf{A}^{\mathrm{o}}$, and compute the inter-agent communication matrix $\mathbf{A}^{\mathrm{a}}_{t}$ at each timestep $t$.
We compute future preferences using Euler integration,
$\mathbf{z}_{i,t+1}=\mathbf{z}_{i,t}+f_{\mathrm{NOD}}\left(\mathbf{z}_{i,t},\mathbf{b}_{t},\mathbf{A}^{\mathrm{a}}_{t}\right)\Delta t,$ (9)
where $\Delta t$ is dataset dependent.
4.3 Communication matrix
We define the communication matrix $\mathbf{A}^{a}_{t}$ as a function of inter-agent proximity. In our baseline BINN model, we define $\mathbf{A}^{a}_{t}$ as the distance between observed positions,
$a^{a}_{(i,j),t}=\lVert\mathbf{p}_{i,t}-\mathbf{p}_{i,t}\rVert^{2}.$ (10)
We also introduce Behavior-Inspired Neural Network+ (BINN+), where we incorporate an informed hypothesis of how agents in the system communicate. For example, for systems in which agents have more influence on each other when they are closer (e.g., human interaction) we define $\mathbf{A}^{a}_{t}$ as the inverse distance between observed positions,
$a^{a}_{(i,j),t}=\frac{1}{\lVert\mathbf{p}_{i,t}-\mathbf{p}_{i,t}\rVert^{2}+\epsilon}$ (11)
where $\epsilon$ is small.
4.4 Decoder
Our decoding network $D_{x}$, is an MPNN [54] that maps agent preferences to predictions of agent states. At every timestep $t$, the latent preferences of the multi-agent system are represented as a fully-connected graph with node values determined by the preferences of each agent.
Formally, our decoder $D_{x}$ performs the following message passing functions for agent $i$ at timestep $t$:
$\quad\quad\hat{\mathbf{x}}^{\prime}_{i,t}$ $=f^{x}_{\mathrm{dec}}\left(\mathbf{z}_{i,t}\right),$ (12)
$v\rightarrow e:$ $\quad\quad\mathbf{m}^{x}_{\left(i,k\right),t}$ $=f_{v\rightarrow e}^{{x}}\left(\hat{\mathbf{x}}^{\prime}_{i,t},\hat{\mathbf{x}}^{\prime}_{k,t}\right),$ (13)
$e\rightarrow v:$ $\quad\quad\hat{\mathbf{x}}_{i,t}$ $=f_{e\rightarrow v}^{{x}}\left(\textstyle\sum_{k\neq i}\mathbf{m}^{x}_{\left(i,k\right),t},\hat{\mathbf{x}}^{\prime}_{i,t}\right),$ (14)
where $f_{\mathrm{dec}}^{{x}}$, $f_{v\rightarrow e}^{x}$, and $f_{{e\rightarrow v}}^{{x}}$ are 3-layer MLPs.
4.5 Loss function
We train our model using the three component loss function,
$\mathcal{L}=\mathcal{L}_{\mathrm{pred}}+\gamma_{1}\mathcal{L}_{\mathrm{recon}}+\gamma_{2}\mathcal{L}_{\mathrm{latent}}$ (15)
where $\mathcal{L}_{\mathrm{pred}}$ is the prediction loss, $\mathcal{L}_{\mathrm{recon}}$ is the reconstruction loss and $\mathcal{L}_{\mathrm{latent}}$ is the latent loss.
The prediction loss, $\mathcal{L}_{\mathrm{pred}}$, is defined as the dissimilarity between the ground truth future state $\mathbf{x}_{i,t}$, and the predicted future state $\hat{\mathbf{x}}_{i,t}$,
$\mathcal{L}_{\mathrm{pred}}=\frac{1}{T\mathcal{N}_{a}}\sum_{t=1}^{T}\sum_{i=1}^{\mathcal{N}_{a}}\Big{\lVert}\mathbf{x}_{i,t}-\hat{\mathbf{x}}_{i,t}\Big{\rVert}^{2},$ (16)
and encourages accuracy of future state prediction. The reconstruction loss, $\mathcal{L}_{\mathrm{recon}}$, is defined as the dissimilarity between the ground truth initial state $\mathbf{x}_{i,0}$, and the reconstructed initial state,
$\mathcal{L}_{\mathrm{recon}}=\frac{1}{\mathcal{N}_{a}}\sum_{i=1}^{\mathcal{N}_{a}}\Big{\lVert}\mathbf{x}_{i,0}-\left(D_{x}\circ E_{z}\right)\left(\mathbf{x}_{i,0}\right)\Big{\rVert}^{2},$ (17)
and encourages the decoder to function as the inverse of the encoder. The latent loss, $\mathcal{L}_{\mathrm{latent}}$, is defined as the dissimilarity of preferences predicted by the dynamical model and those encoded by the preference encoder,
$\mathcal{L}_{\mathrm{latent}}=\frac{1}{T\mathcal{N}_{a}}\sum_{t=1}^{T}\sum_{i=1}^{\mathcal{N}_{a}}\Big{\lVert}\mathbf{z}_{i,t+1}-f_{\mathrm{NOD}}\left(\mathbf{z}_{i,t}\right)\Big{\rVert}^{2},$ (18)
and encourages alignment between the representations encoded by $E_{z}$ and those predicted by $f_{\mathrm{NOD}}$.
5 Experiments
In this section, we highlight the utility of our model for discovering interpretable representations useful for dimensionality reduction, and its efficacy for long-horizon trajectory prediction. Dataset and training details are provided in Appendices A and B.
5.1 Interpretability of the latent space
We demonstrate the interpretability of our latent representations on both the pendulum and double pendulum datasets.
Figure 4: Mutually exclusive pendulum preferences. We show the observed, and latent representations of the pendulum bob. (Top-bottom) The observed position and velocity of the pendulum bob; the learned representation of agent preferences on a 2-dimensional latent space, and the learned representation of agent preferences on a 1-dimensional latent space. In the 2-dimensional space, preferences are perfectly out of phase indicating that they are mutually exclusive.
Pendulum.
We model the pendulum as a single agent system represented as a graph with a single node. We construct a graphical representation at every timestep and set the node feature to the concatenation of the position and velocity of the pendulum bob.
The relationship between observed physical states and learned preferences is shown in Figure 4. Our model suggest the pendulum is a 1-dimensional system which is consistent with physical understanding. We can see this in two ways. First, since preferences $z_{1}$ and $z_{2}$ switch signs when the pendulum bob changes direction, we can interpret the learned categories as a desire to move clockwise ($z_{1}>0$) and a desire to move counter-clockwise ($z_{2}>0$). Moreover, the two preferences are perfectly out-of-phase and therefore mutually exclusive. This is observation is supported by consideration of the learned belief matrix whose entries satisfy the condition for mutual exclusivity, i.e., $a^{o}_{12},a^{o}_{21}\leq 0$,
$\mathbf{A}^{o}=\begin{bmatrix}0&-1.2947\\ -0.9147&0\end{bmatrix}.$ (19)
The conditions for mutual exclusivity are described in Section 3.1.
We empirically verify that the dynamics of this system can be captured with a 1-dimensional latent space. The prediction error for models with a 2-dimensional and 1-dimensional latent space are comparable i.e., $1.9718$e33-3- 3 MSE and $1.9916$e33-3- 3 MSE respectively. Qualitative results are shown in Figure 4 (bottom).
(a) Example trajectories.
(b) Bifurcation diagram.
Figure 5: Pendulum preference bifurcation diagram. (a) We show a pendulum trajectory with clockwise motion (blue) and counterclockwise motion (orange). (b) The initial preferences are near neutral, but large environmental inputs drive preferences in opposite directions resulting in opposite motions in the physical space.
In addition to providing a mechanism for identifying opportunities for dimensionality reduction, the nonlinear opinion dynamics inductive bias also provides a mechanism for controlling preference formation through the environmental input $b_{ij}$. We show trajectories with different rotational directions in Figure 5(a), and their corresponding environmental input latent preference pairs, ($b_{ij}$, $z_{ij}$), on the bifurcation diagram in Figure 5(b). The initial preferences for both trajectories are near neutral, but the initial environmental inputs are far from zero. The negative environmental input drives the initial preference of the blue trajectory toward an equilibrium in the preference space that corresponds to a strongly negative preference, and a clockwise motion in the physical space. The positive environmental input drives the initial preference of the orange trajectory toward an equilibrium in the preference space that corresponds to a strongly positive preference, and a counterclockwise motion in the physical space.
Double pendulum.
We use the double pendulum dataset to demonstrate interpretability in a more complex setting. We model the double pendulum as a two agent system represented as a graph with a two nodes. We construct a graphical representation at every timestep and set node features to the concatenation of the position and velocity of the corresponding double pendulum bob.
Figure 6: Non-mutually exclusive double pendulum preferences. We show the observed, and latent representations of each double pendulum bob. (Top-bottom) The observed position and velocity of the first bob; the learned preferences of the first bob; the observed position and velocity of the second bob; the learned preferences of the second bob. When preference switch, physical bob’s motion switches direction.
The relationship between the observed physical states and the learned preferences is shown in Figure 6. For the first bob, when the preference $z_{1,1}$ switches from greater (less) than to less (greater) than preference $z_{1,2}$, the physical bob switches from clockwise (counterclockwise) to counterclockwise (clockwise) motion. The opposite is true for the second bob. This behavior is observed across examples (see Appendix C.2). In contrast to pendulum preferences, double pendulum preferences are non-mutually exclusive. This observation is supported by the belief matrix, which does not satisfy the mutually-exclusive condition described in Section 3.1,
$\mathbf{A}^{o}=\begin{bmatrix}0&-2.8137\\ 1.1597&0\end{bmatrix}.$ (20)
5.2 Trajectory prediction
We demonstrate the efficacy of our model for trajectory prediction on several datasets. We report the quantitative performance of our model against competitive baselines in Table 1, and show the qualitative performance of our model in Figure 7. Dataset details are provided in Appendix A.
Table 1: Trajectory prediction. Trajectory prediction error on the Mass-Spring, Kuramoto, and TrajNet++ datasets is reported. Our BINN and BINN+ models outperform baseline models on all datasets.
Network Mass-Spring Kuramoto TrajNet++
BINN* $2.88\times 10^{-4}$ $4.52\times 10^{-3}$ $2.16\times 10^{-2}$
BINN+* $\boldsymbol{2.88\times 10^{-4}}$ $\boldsymbol{4.52\times 10^{-3}}$ $\boldsymbol{1.20\times 10^{-2}}$
NRI [19] $3.73\times 10^{-3}$ $8.20\times 10^{-3}$ $8.49\times 10^{-2}$
dNRI [20] $4.86\times 10^{-3}$ $6.41\times 10^{-3}$ $4.00\times 10^{-2}$
Figure 7: Trajectory prediction. We show trajectories predicted by BINN (solid) and the corresponding ground truths (dashed). (Left-right) Examples from mass-spring, Kuramoto oscillator, and TrajNet++ datasets.
5.2.1 Mechanical systems
Mass-spring.
We demonstrate the utility of our model for trajectory prediction on the mass-spring dataset. The system consists of five masses and the dynamics are governed by a second-order linear equation. To model this system, we construct a fully-connected graph with five nodes and initialize node features as the concatenation of the position and velocity of the individual masses.
For both the BINN and BINN+ models we compute the communication matrix using the inter-agent distance defined in Equation (10). Our BINN models outperform both the NRI [19] and dNRI [20] models by an order of magnitude.
Kuramoto oscillator.
The Kuramoto oscillator dataset describes a 5-agent system with dynamics governed by a first-order nonlinear equation. To model this system, we construct a fully-connected graph with five nodes and initialize node features with the vector formed by concatenating the position and velocity of each oscillator.
For both the BINN and BINN+ models we compute the communication matrix using the inter-agent distance defined in Equation (10). On this dataset, our BINN models perform comparably with baseline models.
5.2.2 Pedestrian behavior
TrajNet++ simulator.
We evaluate our model on a synthetic subset of the TrajNet++ [55, 56] dataset where each trajectory contains 5 agents.
For our BINN model, we compute the communication matrix using the inter-agent distance as defined in Equation (10). For our BINN+ model, however, we assume pedestrian sensing scales inversely with inter-agent distance and introduce a prior $\tilde{\mathbf{A}}^{a}_{(i,j),t}$ on the communication matrix,
$\tilde{\mathbf{A}}^{a}_{(i,j),t}=\frac{1}{\lVert\mathbf{p}_{i,t}-\mathbf{p}_{i,t}\rVert^{2}+\epsilon},$ (21)
where $\epsilon$ is very small. We augment $\tilde{\mathbf{A}}^{a}_{(i,j),t}$ with a learned multiplier $\mathbf{A}_{(i,j)}^{\mathrm{pre}}$ which together determine the communication matrix,
Figure 8: TrajNet++ preferences. (Left) We visualize example trajectories from the dataset. The initial position of the trajectory is marked by a blue square and the point of direction change is marked by a black square. (Right) We plot agent preferences for each latent categories. Dominant $z_{i1}$ corresponds to leftwards motion and dominant $z_{i2}$ corresponds to rightward motion. Dominant $z_{i3}$ corresponds to downwards motion and dominant $z_{i4}$ corresponds to upwards motion.
$\mathbf{A}^{a}_{(i,j),t}=\mathbf{A}_{(i,j)}^{\mathrm{pre}}\odot\tilde{\mathbf{A}}^{a}_{(i,j),t}.$ (22)
On this dataset, our BINN and BINN+ models outperform competitive baselines.
The relationship between the observed trajectories and learned preferences is shown in Figure 8. Three consistent trends emerge across examples. First, when the relative magnitudes of $z_{i1}$ and $z_{i1}$ change, the direction of travel changes from right to left. Second, when the relative magnitude of $z_{i3}$ to $z_{i4}$ increases, agents increase their y-velocity. Finally, the dominant $z_{i3}$ in row 1 and 2 results in downwards motion, while the dominant $z_{i4}$ in row 3 results in upwards motion.
6 Conclusion
In this paper we address the challenge of designing a relational inference model with inter-agent relationships that are informed by preferences for multiple categories. To do this, we propose the use of nonlinear opinion dynamics as an inductive bias, and introduce an interpretable and flexible relational inference model. In addition to outperforming competitive existing relational inference models on multiple trajectory prediction datasets, our model provides a mechanisms for identification of mutually exclusive categories, and controlling preference formation.
Limitations and future work.
Similarly to prior work, our BINN models assume the same agents are present throughout a trajectory. Efforts to address this limitation would broaden the applicability of this work.
Appendix
A Datasets
In this section we provide data simulation and post processing steps for the dataset used in Section 5. Trajectories are simulated forward in time with a numerical integrator for a certain timestep and length, all given in Table 3. To reduce the dimensionality of the training data while preserving the dynamics of the system, the simulated trajectories are coarsen according the the frequency given in Table 4 to generate the final training data. We note that since $\Delta t$ is passed into RROD as a fixed hyperparameter for training, the timestep of $1.0$ is arbitrarily chosen for TrajNet++.
A.1 Pendulum
We generate a synthetic dataset of pendulum motion where pendulum dynamics are defined by the second order equation,
$\ddot{\theta}=-\frac{g}{\ell}\sin\theta,$ (23)
with $\ell=1.0$, $g=9.81$, and initial conditions sampled from the uniform distribution, $\theta_{0},\dot{\theta}_{0}\sim\mathcal{U}\left(-0.5\pi,0.5\pi\right)$. The dataset consists of 50000 training trajectories, and 12500 validation and testing trajectories, simulated subject to parameters given in Table 3.
A.2 Double pendulum
We generate a synthetic dataset of double pendulum motion where double pendulum dynamics are defined by the second order equations in [57]. The pendulum arms are defined $\ell_{1}=\ell_{2}=1.0$, pendulum bobs are defined $m_{1}=m_{2}=1.0$, gravity is defined $g=9.81$ and the initial conditions are sampled from the normal distribution, i.e., $\theta_{1,0},\dot{\theta}_{1,0},\theta_{2,0},\dot{\theta}_{2,0}\sim\mathcal{N}\left(0,0.5\right)$. The dataset consists of 50000 training trajectories, and 12500 validation and testing trajectories, simulated subject to parameters given in Table 3.
A.3 Mass-spring
We generate a synthetic dataset of mass spring motion where the mass spring dynamics are defined by the second order linear equation,
$\ddot{\mathbf{r}}_{i}=\sum_{i\neq j}^{\mathcal{N}_{a}}-k_{ij}\left(\mathbf{r}_{i}-\mathbf{r}_{j}\right),$ (24)
where the spring constant $k_{ij}$ is either $2.5$ or 00, and $k_{ij}=k_{ji}$. The initial conditions are sampled from the normal distribution, $\mathbf{r}_{i,0},\dot{\mathbf{r}}_{i,0}\sim\mathcal{N}\left(0,0.3\right)$. The dataset consists of 50000 training trajectories, and 12500 validation and testing trajectories, simulated subject to parameters given in Table 3.
A.4 Kuramoto oscillator
We generate a synthetic dataset of Kuramoto oscillator motion where the oscillator dynamics are defined by the first-order nonlinear equation,
$\dot{\phi}_{i}=\omega_{i}+\sum_{i\neq j}k_{ij}\sin\left(\phi_{i}-\phi_{j}\right),$ (25)
where the coupling constant $k_{ij}$ is either $2.5$ or 00 and $k_{ij}=k_{ji}$. The initial conditions are sampled from the uniform distribution, $\phi_{i,t=0}\sim\mathcal{U}\left(0,2\pi\right)$. The dataset consists of 50000 training trajectories, and 12500 validation and testing trajectories, simulated subject to parameters given in Table 3.
B Training Details
In this section we provide training details on the experiments presented.
All models were trained on Dell Precision 7920 work stations. Each work station contains an Intel Xeon Gold 5220R 24 core CPU, two Nvidia A6000 GPUs, and 256GB of RAM. At training time only 1-GPU is used and the CPU and RAM is shared between two workloads. The hyperparameters used for training is given in Table 4. Furthermore the required memory for training both BINN and baseline models are given in Table 2. The reported values are taken from NVIDIA-smi [58] at the end of the first training epoch, which reports memory usage in terms of mebibyte (MiB) [59]. 1 mebibyte corresponds to 1.04858 megabyte.
Table 2: Memory requirement for RROD, NRI, and dNRI for experiments reported.
Network Mass-Spring Kuramoto TrajNet++
BINN* $496$ MiB $484$ MiB $406$ MiB
NRI $1670$ MiB $1658$ MiB $916$ MiB
dNRI $5008$ MiB $4996$ MiB $2180$ MiB
C Additional Experimental Results
Figure 9: Latent preferences for the double pendulum dataset.
Figure 10: Latent preferences for the double pendulum dataset.
C.1 Separability of Mutually Exclusive Categories
Consider nonlinear opinion dynamics given in Equation (3.1). Given the assumption that $z_{i1}=-cz_{i2}$, we can reduce terms such that belief term
$\sum_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{\mathcal{N}_{\mathrm{o}}}a^{\mathrm{o}}_{jl}z_{il}=a_{21}z_{i2}=-ca_{12}z_{i1}.$ (26)
The communication-belief term
$\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}\sum_{\begin{subarray}{c}l=1\\ l\neq j\end{subarray}}^{\mathcal{N}_{\mathrm{o}}}a^{\mathrm{a}}_{ik}a^{\mathrm{o}}_{jl}z_{kl}$ $=\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}a^{\mathrm{a}}_{ik}a^{\mathrm{o}}_{21}z_{k2},$ (27)
$=\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}-a^{\mathrm{a}}_{ik}ca^{\mathrm{o}}_{12}z_{k1}.$ (28)
And therefore Equation (3.1) can be decoupled into Equation (1), such that
$\dot{z}_{i}=-d_{i}z_{i}+\mathcal{S}\left(u_{i}\left(\tilde{\alpha}_{i}z_{i}+\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{\mathcal{N}_{\mathrm{a}}}\tilde{a}_{ik}z_{k}\right)\right)+b_{i},$
where
$\tilde{\alpha}_{i}$ $=\alpha_{i1}-ca^{o}_{12},$ (30)
$\tilde{a}_{ik}$ $=a^{a}_{i1}-ca^{a}_{i1}a^{0}_{12}.$ (31)
C.2 Double Pendulum
We include additional observation space to preference encoding in Figure 9 and 10.
Table 3: Data generation and post processing parameters for experiments.
Pendulum Double Pendulum Mass-Spring Kuramoto TrajNet++
Integrator RK4 RK4 RK4 RK4 n/a
[…]
Steps $5000$ $5000$ $5000$ $500$ n/a
Coarsening $100$ $100$ $100$ $10$ n/a
Table 4: Hyperparameter for the experiments.
Hyper-parameter Pendulum Double Pendulum Mass-Spring Kuramoto TrajNet++
$\mathcal{N}_{\mathrm{a}}$ $1$ $2$ $5$ $5$ $5$
$\mathcal{N}_{\mathrm{o}}$ $2$ $2$ $4$ $4$ $4$
Epoch $500$ $500$ $1000$ $1000$ $1000$
Batch Size $256$ $256$ $256$ $256$ $256$
Activation ReLU ELU tanh tanh tanh
Optimizer Adam [60] Adam Adam Adam Adam
[…]
Hidden Dimension $64$ $64$ $128$ $128$ $128$
Scheduler N/A N/A StepLR StepLR StepLR
Scheduler Step N/A N/A $200$ $200$ $200$
Scheduler Gamma N/A N/A $0.25$ $0.25$ $0.25$
$\gamma_{1}$ 1.0 1.0 1.0 1.0 1.0
$\gamma_{2}$ 1.0 1.0 1.0 1.0 1.0
Stop spilling…
Stop spilling…